OpenStudy (anonymous):

For the following system, if you isolated x in the first equation to use the Substitution Method, what expression would you substitute into the second equation? 2x + 6y = 18 3x − 2y = 6 −3y + 9 −3y − 9 3y − 9 3y + 9

OpenStudy (welshfella):

2x + 6y = 18 divide through by 2 x + 3y = 9 can you continue/ x = ?

OpenStudy (welshfella):

* subtract 3y from both sides

OpenStudy (welshfella):

are you with me?

OpenStudy (anonymous):

yes

OpenStudy (welshfella):

good ok

OpenStudy (welshfella):

so which one do you think its is

OpenStudy (welshfella):

x + 3y = 9 - 3y -3y